### v3

@@ 3.2 Datasets
The tidal data used in this paper were obtained from ReDAPT project, originating from a tidal energy test site near the Fall of Warness on one of the Orkney Islands. The location can be seen in Figure 6. The TCS data were collected from four-beam Acoustic Doppler Current Profilers instruments deployed on the seabed gravity mooring frames at the tidal energy test site. To comprehensively capture the three-dimensional flow dynamics that determine the magnitude and direction of tidal currents, this study selects the velocity components in the X, Y, and Z directions as forecasting targets. Multi-level forecasting is performed at depths of 22 m, 23 m, and 24 m below sea level. The recordings span 7 days, with data collected at 1 s intervals.
@@
The original data are stored using .mat file format. We first perform preprocessing. Due to malfunctions in reading instruments or interruptions caused by sudden movements near recording sensors, the data contain missing values and outliers. For consecutive missing segments, we substitute a small positive constant on the normalized scale $\varepsilon=0.01$, treating it as a near-zero measurement noise floor to maintain sequence continuity rather than interpreting them as true zeros [33]. One method to remove outliers is to pre-filter high-frequency content in the data [16]. This research uses an adaptive filter to remove high-frequency noise and erroneous data from the dataset. Table 2 shows statistics of the data at different water depths.
@@ 3.3 Evaluation metrics
Figure 7: The stacked bar chart represents the metric values of different models for 10-step forecasting. 22m indicates below sea level 22m.
@@ 3.4 Performance comparisons of different models
We select the following models as baselines: XGBoost [34], LSTM [35], GRU [36], Transformer [37], Reformer [38], iTransformer [39], Autoformer [40], Pyraformer [41], DLinear [42], FEDformer [43] and FiLM [44]. The comparison models’ hyperparameter design refer to the work [45].
@@ 3.5 Performance comparisons of different models
Table 3: Metric values of different models for 30-step forecasting.
@@
XGBoost [34] 0.2546 0.1190 923.65 0.2562 0.1203 899.87 0.2573 0.1212 922.76
LSTM [35] 0.2096 0.1929 979.73 0.2398 0.2147 979.39 0.2140 0.2987 1009.88
GRU [36] 0.2877 0.1359 1452.25 0.2789 0.1283 1330.46 0.2820 0.1327 1349.83
Transformer [37] 0.1359 0.0779 102.20 0.1363 0.0779 130.76 0.1363 0.0743 150.31
Reformer [38] 0.1659 0.1031 120.16 0.1709 0.1095 147.18 0.1710 0.1077 182.59
iTransformer [39] 0.1213 0.0784 109.02 0.1206 0.0786 129.83 0.1193 0.0775 204.86
Autoformer [40] 0.2562 0.2351 257.86 0.2896 0.3117 308.84 0.2235 0.1891 264.44
Pyraformer [41] 0.1306 0.0742 96.33 0.1287 0.0724 118.99 0.1272 0.0719 178.73
DLinear [42] 0.1562 0.1061 124.45 0.1557 0.1063 132.49 0.1538 0.1050 185.70
FEDformer [43] 0.1357 0.0706 122.36 0.1335 0.0701 143.40 0.1341 0.0700 194.76
FiLM [44] 0.1536 0.1108 136.99 0.1528 0.1109 147.70 0.1512 0.1096 211.73
@@
Table 4: Metric values of different models for 60-step forecasting.
BSL 22m 23m 24m
Metrics MAE MSE MAPE (%) MAE MSE MAPE (%) MAE MSE MAPE (%)
XGBoost [34] 0.3584 0.2222 996.64 0.3601 0.2235 1024.54 0.3612 0.2244 966.57
LSTM [35] 0.3327 0.1683 1508.65 0.3582 0.1705 1499.47 0.3503 0.1651 1558.95
GRU [36] 0.3285 0.1792 1791.93 0.3238 0.1808 1508.45 0.3039 0.1659 1620.99
Transformer [37] 0.2129 0.1656 135.84 0.2113 0.1657 176.44 0.2106 0.1645 250.67
Reformer [38] 0.2280 0.1831 156.18 0.2273 0.1840 181.79 0.2261 0.1868 246.09
iTransformer [39] 0.2026 0.1795 166.77 0.2020 0.1804 185.08 0.2000 0.1786 230.24
Autoformer [40] 0.2714 0.2641 265.11 0.2912 0.2911 280.11 0.2622 0.2543 262.56
Pyraformer [41] 0.2068 0.1618 134.66 0.2012 0.1604 156.04 0.2007 0.1596 227.37
DLinear [42] 0.2320 0.1927 165.81 0.2319 0.1937 180.28 0.2294 0.1920 229.17
FEDformer [43] 0.2113 0.1760 174.93 0.2093 0.1767 195.22 0.2072 0.1750 296.24
FiLM [44] 0.2209 0.1997 181.47 0.2205 0.2010 199.22 0.2186 0.1995 255.24
  • 11footnotemark: 1
    Refer to the table 3 for the abbreviations and color codes.
@@
Metrics WECN w/o TPE WECN-RS WECN
MAE 0.081 0.065 0.0245
MSE 0.027 0.0047 0.0029
MAPE (%) 82.9 59.4 24.74
@@
Figure 7 employs a stacked bar chart to illustrate the four evaluation metrics obtained in the 10-step forecast, while Tables 3 and 4 summarize the results for the 30-step and 60-step forecasts, respectively. The reported values represent the averages computed across different directions of tidal current speeds. The WECN achieves a MAE of 0.0250, MSE of 0.0028, and MAPE of 36.23%. Compared to the best-performing model other than WECN, the MAE and MSE decrease by 0.23% and 0.37%, respectively, while MAPE increases by 7.57%. Relative to the second-best model, the reductions in MAE and MSE are 3.47% and 6.66%, respectively, accompanied by a 2.08% increase in MAPE. Overall, WECN achieves the best results in most cases. These findings demonstrate that explicitly modeling the multi-periodic characteristics of tidal currents substantially enhances forecasting accuracy. Moreover, the proposed framework maintains strong performance across datasets collected at different water depths, underscoring its robust generalization capability. We attribute this superior performance to a fundamental architectural advantage. Unlike Transformer-based models, which possess only a weak inductive bias of permutation invariance and must learn temporal ordering from scratch through positional encodings, the WECN leverages the strong inductive biases of locality and translation invariance inherent in its convolutional backbone. The Discrete Wavelet Transform (DWT) module further reinforces these biases by identifying the most salient periodic components within the data. Subsequently, the 1D-to-2D reshaping operation transforms periodicity from a learnable feature into an intrinsic structural property of the input tensor. This effectively reformulates the complex task of modeling long-range temporal dependencies into the simpler objective of capturing short-range spatial dependencies—a domain in which 2D CNNs excel.
Tables 3 and 4 show that as the forecasting time steps increase, the performance of all models significantly decreases. The previous step’s forecasting is the next step’s input in multi-step forecasting. If there is an error in each forecasting step, these errors accumulate with increasing time steps, resulting in increasingly inaccurate forecasting. Despite the forecasting horizon being extended to 30 and 60 steps, WECN effectively mitigates the degradation in predictive accuracy to a certain extent. This robustness can be attributed to WECN’s unique architectural design, which transforms the cascading temporal errors typical of 1D models into a more controllable spatial error diffusion, thereby ensuring stability in multi-step forecasting. Table 3 indicates that in the 30-step forecasting task, compared to the best model other than WECN, MAE increases by 1.74%, MSE decreases by 1.99%, and MAPE decreases by 9.12%. Compared to the second-best model other than WECN, MAE decreases by 4.89%, MSE decreases by 5.49%, and MAPE decreases by 12.20%. Table 4 shows that the model continues to perform well in the 60-step forecast, especially with data at 23m below sea level. Among the three datasets, the data at 23m exhibits the strongest non-stationarity. Multi-periodicity includes non-stationarity within the series. Therefore, the results validate the framework’s effective learning of multi-periodicity.
@@ 3.6.1 Ablation study of the DWT module
To further verify the framework’s ability to capture local periodicity, we added sinusoidal local periodicity to the original data. To control locality, a Gaussian window function was used. The Gaussian window was generated based on the relationship between time, the center offset, and the window width. Then, the Gaussian window was multiplied by a sine wave to create a locally periodic signal, which was added to the original time series data, resulting in a new dataset named $23\mathrm{m}_{\mathrm{new}}$. Next, the WECN and WECN-FFT models are used to forecast both the original and new datasets. Moreover, we optimize the hyperparameters for each model separately. The performance metrics for forecasting are shown in Table 5.
@@ 4 Conclusion
In this research, we developed a multi-periodicity learning framework that is capable of learning local periodicity. We extract dominant periods via DWT and feed them into the predictor without reconstructing components. Then, the TPE algorithm is integrated to optimize hyperparameters, forming the Wavelet-Enhanced Convolutional Network (WECN) framework for tidal current speed forecasting. Numerical results demonstrate that the proposed framework attains a 10-step average MAE of 0.025. Relative to the strongest baseline, error reductions range from 0.23% to 6.66%. In practical applications, tidal current time series exhibit multi-periodicity arising from the interaction of different tidal components and local flow conditions. The proposed framework models these multi-periodic features, enabling adaptation to the complex and dynamic patterns in real marine environments, which holds significant importance for practical engineering applications. Regarding generalization ability, the framework demonstrates applicability across different datasets and can be used for similar multi-periodicity time series tasks. The results indicate that the framework can provide valuable support for integrating tidal energy into the electrical grid.
Despite its promising performance, the proposed framework still faces certain limitations that need to be addressed in future research. The current model uses tidal current time series for forecasting without considering additional features. Future work will aim to improve prediction accuracy by integrating knowledge-driven approaches.

### v4

@@ 3.2 Datasets
The tidal data used in this paper were obtained from ReDAPT project, originating from a tidal energy test site near the Fall of Warness on one of the Orkney Islands. The location can be seen in Figure 6. The TCS data were collected from four-beam Acoustic Doppler Current Profilers instruments deployed on the seabed gravity mooring frames at the tidal energy test site.
The dataset was recorded continuously over a seven-day period with a high-resolution sampling interval of 1 s, allowing for detailed temporal analysis. The 1 s sampling interval is particularly important for capturing fine-grained dynamics and turbulent fluctuations inherent in the system.
The dataset is deliberately designed to support an evaluation of model generalization under distributional variations. Specifically, the velocity components along the X, Y, and Z directions are treated as independent forecasting targets to capture direction-dependent tidal flow dynamics. In addition, multi-level forecasting is conducted at three different depths (22,m, 23,m, and 24,m below sea level), introducing hydrodynamic variations induced by depth differences. These settings expose the model to multiple related but non-identical data distributions, enabling an assessment of whether the learned representations generalize beyond a single channel or a single depth. This design allows the proposed method to be evaluated across multiple velocity components and depths within a consistent physical environment, providing a controlled yet non-trivial test of generalization under limited but realistic high-frequency tidal measurements.
@@
The original data are stored using .mat file format. We first perform preprocessing. Due to malfunctions in reading instruments or interruptions caused by sudden movements near recording sensors, the data contain missing values and outliers. For non-consecutive missing values, gaps are filled using the average of neighboring time points. This approach provides a reasonable local estimate while preserving the temporal continuity of the series without introducing significant bias. For consecutive missing segments, the data are directly replaced with a default value of 0.01. This small constant is used to ensure numerical stability and consistency during processing, to clearly distinguish imputed values from true zeros, and to minimize their impact on the overall statistical properties of the dataset. [Chavat2022ECDUY]. One method to remove outliers is to pre-filter high-frequency content in the data [5589321]. This research uses an adaptive filter to remove high-frequency noise and erroneous data from the dataset. Table 2 shows statistics of the data at different water depths.
@@ 3.4 Performance comparisons of different models
We select the following models as baselines: XGBoost [chen2016xgboost], LSTM [hochreiter1997long], GRU [chung2014empirical], Reformer [Kitaev2020Reformer], Autoformer [wu2021autoformer], DLinear [Zeng_Chen_Zhang_Xu_2023], FEDformer [pmlr-v162-zhou22g] and FiLM [NEURIPS2022_524ef58c]. The comparison models’ hyperparameter design refer to the work [wang2024tssurvey].
@@ 3.5 Performance comparisons of different models
Table 3: Metric values of different models for multi-step forecasting.
(a) 10-step forecasting
@@
XGBoost 0.2489 0.1137 485.51 0.2504 0.1150 507.83 0.2517 0.1160 505.37
LSTM 0.1014 0.0210 335.15 0.0926 0.0187 259.02 0.1025 0.0223 297.83
GRU 0.2530 0.1001 1242.34 0.2666 0.1090 1154.99 0.2589 0.1072 1154.47
Reformer 0.0662 0.0100 47.37 0.0646 0.0103 62.00 0.0662 0.0107 99.90
Autoformer 0.1601 0.0891 176.06 0.1961 0.1348 236.62 0.1699 0.1060 297.15
DLinear 0.0347 0.0062 35.17 0.0344 0.0062 41.87 0.0340 0.0061 62.68
FEDformer 0.0518 0.0052 50.68 0.0429 0.0037 40.91 0.0448 0.0039 164.95
FiLM 0.0257 0.0033 29.63 0.0255 0.0033 29.15 0.0251 0.0032 47.68
WECN (Ours) 0.0254 0.0025 30.14 0.0245 0.0029 24.74 0.0253 0.0030 53.83
(b) 30-step forecasting
BSL 22m 23m 24m
Metrics MAE MSE MAPE (%) MAE MSE MAPE (%) MAE MSE MAPE (%)
XGBoost [chen2016xgboost] 0.2546 0.1190 923.65 0.2562 0.1203 899.87 0.2573 0.1212 922.76
LSTM [hochreiter1997long] 0.2096 0.1929 979.73 0.2398 0.2147 979.39 0.2140 0.2987 1009.88
GRU [chung2014empirical] 0.2877 0.1359 1452.25 0.2789 0.1283 1330.46 0.2820 0.1327 1349.83
Reformer [Kitaev2020Reformer] 0.1659 0.1031 120.16 0.1709 0.1095 147.18 0.1710 0.1077 182.59
Autoformer [wu2021autoformer] 0.2562 0.2351 257.86 0.2896 0.3117 308.84 0.2235 0.1891 264.44
DLinear [Zeng_Chen_Zhang_Xu_2023] 0.1562 0.1061 124.45 0.1557 0.1063 132.49 0.1538 0.1050 185.70
FEDformer [pmlr-v162-zhou22g] 0.1357 0.0706 122.36 0.1335 0.0701 143.40 0.1341 0.0700 194.76
FiLM [NEURIPS2022_524ef58c] 0.1536 0.1108 136.99 0.1528 0.1109 147.70 0.1512 0.1096 211.73
@@
(c) 60-step forecasting
BSL 22m 23m 24m
Metrics MAE MSE MAPE (%) MAE MSE MAPE (%) MAE MSE MAPE (%)
XGBoost [chen2016xgboost] 0.3584 0.2222 996.64 0.3601 0.2235 1024.54 0.3612 0.2244 966.57
LSTM [hochreiter1997long] 0.3327 0.1683 1508.65 0.3582 0.1705 1499.47 0.3503 0.1651 1558.95
GRU [chung2014empirical] 0.3285 0.1792 1791.93 0.3238 0.1808 1508.45 0.3039 0.1659 1620.99
Reformer [Kitaev2020Reformer] 0.2280 0.1831 156.18 0.2273 0.1840 181.79 0.2261 0.1868 246.09
Autoformer [wu2021autoformer] 0.2714 0.2641 265.11 0.2912 0.2911 280.11 0.2622 0.2543 262.56
DLinear [Zeng_Chen_Zhang_Xu_2023] 0.2320 0.1927 165.81 0.2319 0.1937 180.28 0.2294 0.1920 229.17
FEDformer [pmlr-v162-zhou22g] 0.2113 0.1760 174.93 0.2093 0.1767 195.22 0.2072 0.1750 296.24
FiLM [NEURIPS2022_524ef58c] 0.2209 0.1997 181.47 0.2205 0.2010 199.22 0.2186 0.1995 255.24
@@
Model MAE MSE MAPE (%)
FiLM w/o TPE 0.0255 0.0033 29.15
FiLM 0.0248 0.0029 27.66
FEDformer w/o TPE 0.0429 0.0037 40.91
FEDformer 0.0315 0.0032 31.20
WECN w/o TPE 0.0810 0.0270 82.90
WECN-RS 0.0650 0.0047 59.40
WECN 0.0245 0.0029 24.74
@@
Table 3(a) presents the four evaluation metrics obtained under the 10-step forecasting setting using a stacked bar chart, while Table 3(b) and Table 3(c) summarize the results for the 30-step and 60-step forecasts, respectively.
Table 3(b) and Table 3(c) show that the reported values are averaged across different directions of tidal current speeds. Under the 10-step forecasting setting, WECN achieves an MAE of 0.0251, an MSE of 0.0028, and a MAPE of 36.24%. Compared with the best-performing model other than WECN, the MAE and MSE decrease by 1.44% and 14.29%, respectively, while the MAPE increases by 2.11%. Relative to the second-best model, the MAE and MSE are reduced by 27.06% and 54.59%, respectively, together with a 22.19% reduction in MAPE. Overall, WECN achieves the best results in most cases, indicating that explicitly modeling the multi-periodic characteristics of tidal currents improves forecasting accuracy. In addition, the proposed framework maintains consistent performance across datasets collected at different water depths, demonstrating robust generalization ability.
Table 3(b) and Table 3(c) further indicate that forecasting performance degrades as the prediction horizon increases. In recursive multi-step forecasting, the prediction at the previous step is used as the input for the next step. Consequently, step-wise errors accumulate as the forecasting horizon grows, leading to reduced accuracy. Even when the horizon is extended to 30 and 60 steps, WECN alleviates performance degradation to a certain extent. This behavior is attributed to the architectural design of WECN, which converts the cascading temporal error accumulation common in one-dimensional models into a more controllable spatial error diffusion process, thereby improving stability in multi-step forecasting. As shown in Table 3(b), for the 30-step forecasting task, compared with the best model other than WECN, MAE, MSE, and MAPE decrease by 8.88%, 1.99%, and 24.59%, respectively. Relative to the second-best model, the reductions reach 21.09% in MAE, 34.94% in MSE, and 21.54% in MAPE. Table 3(c) shows that the proposed model continues to perform well in the 60-step forecasting task, particularly for the dataset collected at 23 m below sea level. Among the three datasets, the 23 m dataset exhibits the strongest non-stationarity. Since multi-periodicity includes non-stationary components within the time series, these results confirm that the proposed framework effectively learns multi-periodic characteristics.
@@ 3.6.1 Ablation study of the DWT module
To further verify the framework’s ability to capture local periodicity, we added sinusoidal local periodicity to the original data. To control locality, a Gaussian window function was used. The Gaussian window was generated based on the relationship between time, the center offset, and the window width. Then, the Gaussian window was multiplied by a sine wave to create a locally periodic signal, which was added to the original time series data, resulting in a new dataset named $23\mathrm{m}_{\mathrm{new}}$. Next, the WECN and WECN-FFT models are used to forecast both the original and new datasets. Moreover, we optimize the hyperparameters. The performance metrics for forecasting are shown in Table 4.
@@
Results can be explained by the mismatch between global Fourier bases and non-stationary tidal dynamics. Real tidal-current sequences often contain quasi-periodic components with time-varying amplitude and phase, so the energy of a modulated sinusoid is not concentrated at a single discrete FFT bin but spreads to nearby frequencies due to finite-length truncation and windowing, which makes the extracted dominant periods sensitive to the chosen window and noise. By contrast, DWT implements a multiresolution filter bank that partitions the spectrum into dyadic subbands, and the periodicity strength can be measured by the subband coefficient energy. This representation is stable under mild frequency drift because the energy remains within the same subband even when the instantaneous frequency moves inside that band. Therefore, when locally windowed sinusoids are injected into $23\mathrm{m}_{\mathrm{new}}$, DWT yields a clearer separation between long-term tidal trends and short-term oscillatory bursts, whereas FFT-based features are more affected by spectral leakage and band spreading, which degrades WECN-FFT and improves WECN in the reported comparisons.
@@ 3.6.2 Ablation study of the TPE module
To avoid restricting the analysis to the proposed model, the same TPE optimization procedure is also applied to two strong baseline models in this setting, namely FiLM and FEDformer, which show the second- and third-best performance among the compared baselines in Table 5.
@@
4 Discussion
4.1 Practical implications
Accurate short-term tidal current speed forecasting serves as an operational module for tidal energy systems and grid integration. The proposed WECN models multi-periodicity arising from the interaction between tidal constituents and local flow conditions. This modelling design directly targets engineering deployment scenarios related to tidal energy integration into electrical grids. From a deployment perspective, the framework produces multi-step forecasts, which support operational decisions in cases where single-step prediction does not provide sufficient information for system controllability and reliability.
The study adopts a recursive multi-step forecasting strategy to generate a forecast trajectory. This trajectory supports:
  • short-horizon control actions, including set-point adjustment and power smoothing;
  • longer-horizon operational planning, including scheduling and risk-aware decision-making based on an evolving future sequence rather than a single point estimate.
4.2 Limitations
Several limitations are associated with the current design. First, the current framework relies exclusively on tidal current time series and does not incorporate exogenous variables or knowledge-driven constraints. Accordingly, the integration of knowledge-driven approaches is identified as a direction for future work. In addition, the data processing pipeline applies missing-value imputation and predefined filtering operations, and hyperparameter selection is performed under a bounded evaluation budget. These design choices influence robustness and require further examination under varying data quality conditions and operational constraints.
@@ 5 Conclusion
In this research, we developed a multi-periodicity learning framework that is capable of learning local periodicity. We extract dominant periods via DWT and feed them into the predictor without reconstructing components. Then, the TPE algorithm is integrated to optimize hyperparameters, forming the Wavelet-Enhanced Convolutional Network (WECN) framework for tidal current speed forecasting. Numerical results demonstrate that the proposed framework attains a 10-step average MAE of 0.025. Relative to the strongest baseline, error reductions range from 0.23% to 6.66%. In practical applications, tidal current time series exhibit multi-periodicity arising from the interaction of different tidal components and local flow conditions. The proposed framework models these multi-periodic features, enabling adaptation to the complex and dynamic patterns in real marine environments, which holds significant importance for practical engineering applications. Regarding generalization ability, the framework demonstrates applicability across different datasets and can be used for similar multi-periodicity time series tasks. The results indicate that the framework can provide valuable support for integrating tidal energy into the electrical grid. The results indicate that the framework can provide effective support for integrating tidal energy into the electrical grid. However, the current framework relies solely on tidal current time series, indicating the need to incorporate additional information and constraints in future work to extend the method.